Find the limit. 5.3 + 9 lim 100 20.r3 3.x2 + 9
Answer #1

Sool L = lim 5x3+g dox ²-3x²+g L = clim x8 [5+ +973) Roo X [20-24 + 2 2 2 L = 5+o 20-oto សា L = 5 [ L = 4 JA